Find the value of b for which the equation
(1 )( ye^2xy + x )dx + bxe^2xy dy = 0
is exact, and hence solve it for that value of b
1
Expert's answer
2020-03-26T11:01:53-0400

"P(x,y)=bxe^{2xy};Q(x,y)=ye^{2xy}+x"

"P_x=be^{2xy}+2bye^{2xy}"

"Q_y=e^{2xy}+2xye^{2xy}"

The equation is exact if "P_x=Q_y", hence

"be^{2xy}(1+2xy)=e^{2xy}(1+2xy)"

"b=1"

The solution of the equation "R(x,y)=const" , where

"R=\int Pdy=\int xe^{2xy}dy=\frac{e^{2xy}}{2}+h(x)"

"R_x=Q" hence

"ye^{2xy}+h'=ye^{2xy}+x"

"h'=x"

"h=\frac{x^2}{2}"

The solution of the equation

"\frac{1}{2}(e^{2xy}+x^2)=const"